# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Abstract
The performance of a model trained with noisy labels is often improved by simply retraining the model with its own predicted hard labels (i.e., $1$/00 labels). Yet, a detailed theoretical characterization of this phenomenon is lacking. In this paper, we theoretically analyze retraining in a linearly separable setting with randomly corrupted labels given to us and prove that retraining can improve the population accuracy obtained by initially training with the given (noisy) labels. To the best of our knowledge, this is the first such theoretical result. Retraining finds application in improving training with label differential privacy (DP) which involves training with noisy labels. We empirically show that retraining selectively on the samples for which the predicted label matches the given label significantly improves label DP training at no extra privacy cost; we call this consensus-based retraining. For e.g., when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining.
@@ 1 Introduction
We study the simple idea of retraining an already trained model with its own predicted hard labels (i.e., $1$/00 labels and not the raw probabilities) when the given labels with which the model is initially trained are noisy. This is a simple yet effective way to boost a model’s performance in the presence of noisy labels. More formally, suppose we train a discriminative model $\mathcal{M}$ (for a classification problem) on a dataset of $n$ samples and noisy label pairs $\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j=1}^{n}$. Let $\widehat{\bm{\theta}}_{0}$ be the final learned weight/checkpoint of $\mathcal{M}$ and let $\widetilde{\mathrm{y}}_{j}=\mathcal{M}(\widehat{\bm{\theta}}_{0},{\mathbf{x}}_{j})$ be the current checkpoint’s predicted hard label for sample ${\mathbf{x}}_{j}$. Now, we propose to retrain $\mathcal{M}$ with the $\widetilde{\mathrm{y}}_{j}$’s in one of two ways.
(i) Full retraining: Retrain $\mathcal{M}$ with $\{({\mathbf{x}}_{j},\widetilde{\mathrm{y}}_{j})\}_{j=1}^{n}$, i.e., retrain $\mathcal{M}$ with the predicted labels of all the samples.
(ii) Consensus-based retraining: Define $\mathcal{S}_{\text{cons}}:=\{j\in\{1,\ldots,n\}\text{ }|\text{ }\widetilde{\mathrm{y}}_{j}=\widehat{\mathrm{y}}_{j}\}$ to be the set of samples for which the predicted label matches the given noisy label; we call this the consensus set. Retrain $\mathcal{M}$ with $\{({\mathbf{x}}_{j},\widetilde{\mathrm{y}}_{j})\}_{j\in\mathcal{S}_{\text{cons}}}$, i.e., retrain $\mathcal{M}$ with the predicted labels of only the consensus set.
@@
The motivation for consensus-based retraining is that matching the predicted and given labels can potentially yield a smaller but much more accurate subset compared to the entire set; such a filtering effect can further improve the model’s performance. As we show in Section 5 (see Tables 3, 4 and 6), this intuition bears out in practice.
(a) Large separation: predicted labels pretty
accurate.
(b) Small separation: predicted labels not as
accurate.
@@
The primary reason for our interest in retraining is that it turned out to be a simple yet effective way to improve training with label differential privacy (DP) whose goal is to safeguard the privacy of labels in a supervised ML problem by injecting label noise (see Section 3 for a formal definition). Label DP is used in scenarios where only the labels are considered sensitive, e.g., advertising, recommendation systems, etc. (Ghazi et al.,, 2021). Importantly, retraining can be applied on top of any label DP training algorithm at no extra privacy cost. Our main algorithmic contribution is empirically demonstrating the efficacy of consensus-based retraining in improving label DP training (Section 5). Three things are worth clarifying here. First, as a meta-idea, retraining is not particularly new; however, its application – especially with consensus-based filtering – as a light-weight way to improve label DP training at no extra privacy cost is new to our knowledge. Second, we are not advocating consensus-based retraining as a SOTA general-purpose algorithm for learning with noisy labels. Third, we do not view full retraining to be an algorithmic contribution; we consider it for theoretical analysis and as a baseline for consensus-based retraining.
@@
(a) In Section 4, we consider a linearly separable binary classification problem wherein the data (feature) dimension is $d$, and we are given randomly flipped labels with the label flipping probability being $p<\frac{1}{2}$ independently for each sample. Our main result is proving that full retraining with the predicted hard labels improves the population accuracy obtained by initially training with the given labels, provided the degree of separation between the classes is large enough and the number of samples $n\in\Big{(}\Theta\Big{(}\frac{d\log d}{(1-2p)^{2}}\Big{)},\Theta\Big{(}\frac{d^{2}}{(1-2p)^{2}}\Big{)}\Big{)}$; see Remark 4 for details. Our results also show that retraining becomes more beneficial as the amount of label noise (i.e., $p$) increases (Remark 4) or as the degree of separation between the classes increases (Remark 5). To the best our knowledge, these are the first theoretical results quantifying the benefits of retraining with predicted hard labels in the presence of label noise.
(b) In Section 5, we show the promise of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) as a simple way to improve the performance of any label DP algorithm, at no extra privacy cost. For e.g., when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining (see Table 2). The corresponding improvement for a small BERT model trained on AG News Subset (a news classification dataset) with $\epsilon=0.5$ label DP is $11.7\%$ (see Table 5).
@@ 2 Related Work
Self-Training (ST). Retraining is similar in spirit to ST (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013; Sohn et al.,, 2020) which is the process of progressively training a model with its own predicted hard labels in the semi-supervised setting. Our main focus is the fully supervised setting; [^1] It is worth mentioning here that we analyze retraining in a semi-supervised setting due to certain technical complications (because of independence not holding) in the supervised setting. But we believe that the crux of our theoretical insights should carry over to the supervised setting. in fact, our proposed consensus-based retraining scheme crucially relies on the given labels for sample selection. This is different from ST (in the semi-supervised setting) which typically selects samples based on the model’s confidence and hence, we call our algorithmic idea of interest retraining to distinguish it from ST. In fact, we show that our consensus-based sample selection strategy leads to better performance than confidence-based sample selection in Appendix I. There is a vast body of work on ST and related ideas; see Amini et al., (2022) for a survey. On the theoretical side also, there are several papers showing and quantifying different kinds of benefits of ST and related ideas (Carmon et al.,, 2019; Raghunathan et al.,, 2020; Kumar et al.,, 2020; Chen et al.,, 2020; Oymak and Gulcu,, 2020; Wei et al.,, 2020; Zhang et al.,, 2022). But none of these works characterize the pros/cons of ST or any related algorithm in the presence of noisy labels. In contrast, we show that retraining can provably improve accuracy in the presence of label noise in Section 4. Empirically, ST-based ideas have been proposed to improve learning with noisy labels (Reed et al.,, 2014; Tanaka et al.,, 2018; Han et al.,, 2019; Nguyen et al.,, 2019; Li et al.,, 2020; Goel et al.,, 2022); but these works do not have rigorous theory.
@@
Label Differential Privacy (DP). Label DP (described in detail in Section 3) is a relaxation of full-data DP wherein the privacy of only the labels (and not the features) is safeguarded (Chaudhuri and Hsu,, 2011; Beimel et al.,, 2013; Wang and Xu,, 2019; Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021; Ghazi et al.,, 2022; Badanidiyuru et al.,, 2023). In this work, we are not trying to propose a SOTA label DP algorithm (with an ingenious noise-injection scheme); instead, we advocate retraining as a simple post-processing step that can be applied on top of any label DP algorithm (regardless of the noise-injection scheme) to improve its performance, at no extra privacy cost. Similar to our goal, Tang et al., (2022) apply techniques from unsupervised and semi-supervised learning to improve label DP training. In particular, one of their steps involves keeping the given noisy label of a sample only if it matches a pseudo-label generated by unsupervised learning. This is similar in spirit to our consensus-based retraining scheme but a crucial difference is that we do not perform any unsupervised learning; we show that matching the given noisy label to the model’s own predicted label is itself pretty effective. Further, unlike work, Tang et al., (2022) do not have any rigorous theory.
@@ 3 Preliminaries
Notation. We use $\Theta(\cdot)$ and $\Omega(\cdot)$ to denote the standard big-$\Theta$ and big-$\Omega$ notations. Specifically, $f(d)=\Theta(g(d))$ if there are absolute constants $c_{1},c_{2}>0$ and a natural number $d_{0}$ such that $c_{1}g(d)\leq f(d)\leq c_{2}g(d)$ for all $d\geq d_{0}$. In addition, $f(d)=\Omega(g(d))$ if there is an absolute constant $c$ and a natural number $d_{0}$ such that $f(d)\geq cg(d)$ for all $d\geq d_{0}$. Throughout the paper, we sometimes use the big-$\Theta$ notation to absorb different absolute constants. For any positive integer $m\geq 1$, we denote the set $\{1,\ldots,m\}$ by $[m]$. For any vector ${\mathbf{v}}$, we denote its $i^{\text{th}}$ coordinate by $\mathrm{v}^{(i)}$. Let ${\mathbf{e}}_{i}$ denote the $i^{\text{th}}$ canonical vector, i.e., the vector of all zeros except a one in the $i^{\text{th}}$ coordinate. We denote the $\ell_{2}$ norm of a vector ${\mathbf{v}}$ by $\|{\mathbf{v}}\|$, and the operator norm of a matrix ${\mathbf{M}}$ by $\|{\mathbf{M}}\|$. The unit $d$-dimensional sphere (i.e., the set of $d$-dimensional vectors with unit norm) is denoted by $S^{d-1}$. A random variable ${\mathrm{X}}$ is said to be sub-Gaussian with parameter $\sigma^{2}$ (abbreviated as SG($\sigma^{2}$)) if for all $t\in\mathbb{R}$, it holds that $\mathbb{E}\big{[}e^{t({\mathrm{X}}-\mathbb{E}[{\mathrm{X}}])}\big{]}\leq\exp\big{(}\frac{t^{2}\sigma^{2}}{2}\big{)}$. We denote the CDF and complementary CDF (CCDF) of a standard normal variable (i.e., distributed as $\mathcal{N}(0,1)$) by $\Phi(.)$ and $\Phi^{\text{c}}(.)$, respectively.
Definition 1 (Label Differential Privacy (DP)).
@@
Label DP training involves injecting noise into the labels and then training with these noisy labels. The simplest way of injecting label noise to ensure label DP is randomized response (RR) introduced by Warner, (1965). Specifically, suppose we require $\epsilon$-labelDP for a problem with $C$ classes, then the distribution of the output $\widehat{\mathrm{y}}$ of RR when the true label is ${\mathrm{y}}$ is as follows:
$\mathbb{P}\big{(}\hat{\mathrm{y}}={\mathrm{z}}\big{)}=\begin{cases}\frac{e^{\epsilon}}{e^{\epsilon}+C-1}\quad\text{ for }{\mathrm{z}}={\mathrm{y}},\\ \frac{1}{e^{\epsilon}+C-1}\quad\text{ otherwise}.\end{cases}$
Our label noise model in Section 4 (eq. 1) is actually RR for 2 classes. Based on vanilla RR, more sophisticated ways to inject label noise for better performance under label DP have been proposed (Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021). Our empirical results in Section 5 are with RR and the method of Ghazi et al., (2021).
@@ 4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
Here we will analyze full retraining (as introduced in Section 1) for a linear setting with noisy labels. Since full retraining is the only kind of retraining we consider in this section, we will omit the word “full” subsequently in this section.

### v2

@@ Abstract
The performance of a model trained with noisy labels is often improved by simply retraining the model with its own predicted hard labels (i.e., $1$/00 labels). Yet, a detailed theoretical characterization of this phenomenon is lacking. In this paper, we theoretically analyze retraining in a linearly separable setting with randomly corrupted labels given to us and prove that retraining can improve the population accuracy obtained by initially training with the given (noisy) labels. To the best of our knowledge, this is the first such theoretical result. Retraining finds application in improving training with local label differential privacy (DP) which involves training with noisy labels. We empirically show that retraining selectively on the samples for which the predicted label matches the given label significantly improves label DP training at no extra privacy cost; we call this consensus-based retraining. As an example, when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining.
@@ 1 Introduction
We study the simple idea of retraining an already trained model with its own predicted hard labels (i.e., $1$/00 labels and not the raw probabilities) when the given labels with which the model is initially trained are noisy. This is a simple yet effective way to boost a model’s performance in the presence of noisy labels. More formally, suppose we train a discriminative model $\mathcal{M}$ (for a classification problem) on a dataset of $n$ samples and noisy label pairs $\{(\bm{x}_{j},\widehat{y}_{j})\}_{j=1}^{n}$. Let ${\widehat{\bm{\theta}}}_{0}$ be the final learned weight/checkpoint of $\mathcal{M}$ and let $\widetilde{y}_{j}=\mathcal{M}({\widehat{\bm{\theta}}}_{0},\bm{x}_{j})$ be the current checkpoint’s predicted hard label for sample $\bm{x}_{j}$. Now, we propose to retrain $\mathcal{M}$ with the $\widetilde{y}_{j}$’s in one of the following two ways:
  • $(i)$
    Full retraining: Retrain $\mathcal{M}$ with $\{(\bm{x}_{j},\widetilde{y}_{j})\}_{j=1}^{n}$, i.e., retrain $\mathcal{M}$ with the predicted labels of all the samples.
  • $(ii)$
    Consensus-based retraining: Define $\mathcal{S}_{\text{cons}}:=\{j\in\{1,\ldots,n\}\text{ }|\text{ }\widetilde{y}_{j}=\widehat{y}_{j}\}$ to be the set of samples for which the predicted label matches the given noisy label; we call this the consensus set. Retrain $\mathcal{M}$ with $\{(\bm{x}_{j},\widetilde{y}_{j})\}_{j\in\mathcal{S}_{\text{cons}}}$, i.e., retrain $\mathcal{M}$ with the predicted labels of only the consensus set.
@@
The motivation for consensus-based retraining is that matching the predicted and given labels can potentially yield a smaller but much more accurate subset compared to the entire set; such a filtering effect can further improve the model’s performance. As we show in Section 5 (see Tables 3 and 7), this intuition bears out in practice.
(a) Large separation:
predicted labels pretty accurate.
(b) Small separation:
predicted labels not as accurate.
@@
The primary reason for our interest in retraining is that it turned out to be a simple yet effective way to improve training with local [^1] In this paper, we focus only on local label DP. So throughout the paper, we will mostly omit the word “local” for conciseness. label differential privacy (DP) whose goal is to safeguard the privacy of labels in a supervised ML problem by injecting label noise (see Section 3 for a formal definition). Label DP is used in scenarios where only the labels are considered sensitive, e.g., advertising, recommendation systems, etc. (Ghazi et al.,, 2021). Importantly, retraining can be applied on top of any label DP training algorithm at no extra privacy cost. Our main algorithmic contribution is empirically demonstrating the efficacy of consensus-based retraining in improving label DP training (Section 5). Three things are worth clarifying here. First, as a meta-idea, retraining is not particularly new; however, its application – especially with consensus-based filtering – as a light-weight way to improve label DP training at no extra privacy cost is new to our knowledge. Second, we are not advocating consensus-based retraining as a SOTA general-purpose algorithm for learning with noisy labels. Third, we do not view full retraining to be an algorithmic contribution; we consider it for theoretical analysis and as a baseline for consensus-based retraining.
@@
  • In Section 4, we consider a linearly separable binary classification problem wherein the data (feature) dimension is $d$, and we are given randomly flipped labels with the label flipping probability being $p<\frac{1}{2}$ independently for each sample. Our main result is proving that full retraining with the predicted hard labels improves the population accuracy obtained by initially training with the given labels, provided that $p$ is close enough to $\frac{1}{2}$ and the dataset size $n$ satisfies $\frac{d\log n}{(1-2p)}\lesssim n\lesssim\frac{d^{2}}{(1-2p)^{2}}$ ($\lesssim$” means “bounded asymptotically”, ignoring constant factor multiples); see Remark 4.10 for details. In addition, our results show that retraining becomes more beneficial as the amount of label noise (i.e., $p$) increases or as the degree of separation between the classes increases. To the best of our knowledge, these are the first theoretical results quantifying the benefits of retraining with predicted hard labels in the presence of label noise. The analysis of retraining is particularly challenging because of the dependence of the predicted labels on the entire training set and the non-uniform/sample-dependent nature of label noise in the predicted labels; see the discussion after the statement of Theorem 4.8.
  • In Section 5, we show the promise of consensus-based retraining (i.e., retraining on only those samples for which the predicted label matches the given noisy label) as a simple way to improve the performance of any label DP algorithm, at no extra privacy cost. As an example, when training ResNet-18 on CIFAR-100 with $\epsilon=3$ label DP, we obtain $6.4\%$ improvement in accuracy with consensus-based retraining (see Table 2). The corresponding improvement for a small BERT model trained on AG News Subset (a news classification dataset) with $\epsilon=0.5$ label DP is $11.7\%$ (see Table 6).
@@ 2 Related Work
Self-Training (ST). Retraining is similar in spirit to ST (Scudder,, 1965; Yarowsky,, 1995; Lee et al.,, 2013; Sohn et al.,, 2020) which is the process of progressively training a model with its own predicted hard labels in the semi-supervised setting. Our focus in this work is on the fully supervised setting. This is different from ST (in the semi-supervised setting) which typically selects samples based on the model’s confidence and hence, we call our algorithmic idea of interest retraining to distinguish it from ST. In fact, we show that our consensus-based sample selection strategy leads to better performance than confidence-based sample selection in Appendix H. There is a vast body of work on ST and related ideas; see Amini et al., (2022) for a survey. On the theoretical side also, there are several papers showing and quantifying different kinds of benefits of ST and related ideas (Carmon et al.,, 2019; Raghunathan et al.,, 2020; Kumar et al.,, 2020; Chen et al.,, 2020; Oymak and Gulcu,, 2020; Wei et al.,, 2020; Zhang et al.,, 2022). But none of these works characterize the pros/cons of ST or any related algorithm in the presence of noisy labels. In contrast, we show that retraining can provably improve accuracy in the presence of label noise in Section 4. Empirically, ST-based ideas have been proposed to improve learning with noisy labels (Reed et al.,, 2014; Tanaka et al.,, 2018; Han et al.,, 2019; Nguyen et al.,, 2019; Li et al.,, 2020; Goel et al.,, 2022); but these works do not have rigorous theory. In the context of theory on label noise and model’s confidence, Zheng et al., (2020) show that if the model’s predicted score for the observed label is small, then the observed label is likely not equal to the true label. Note that the results of Zheng et al., (2020) pertain to the correctness of the observed labels, whereas our results pertain to the correctness of the predicted labels.
@@
Label Differential Privacy (DP). Label DP (described in detail in Section 3) is a relaxation of full-data DP wherein the privacy of only the labels (and not the features) is safeguarded (Chaudhuri and Hsu,, 2011; Beimel et al.,, 2013; Wang and Xu,, 2019; Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021; Ghazi et al.,, 2022; Badanidiyuru et al.,, 2023). In this work, we are not trying to propose a SOTA label DP algorithm (with an ingenious noise-injection scheme); instead, we advocate retraining as a simple post-processing step that can be applied on top of any label DP algorithm (regardless of the noise-injection scheme) to improve its performance, at no extra privacy cost. Similar to our goal, Tang et al., (2022) apply techniques from unsupervised and semi-supervised learning to improve label DP training. In particular, one of their steps involves keeping the given noisy label of a sample only if it matches a pseudo-label generated by unsupervised learning. This is similar in spirit to our consensus-based retraining scheme but a crucial difference is that we do not perform any unsupervised learning; we show that matching the given noisy label to the model’s own predicted label is itself pretty effective. Further, unlike our work, Tang et al., (2022) do not have any rigorous theory.
@@ 3 Preliminaries
Notation. For two functions $f(n)$ and $g(n)$, we write $f(n)\lesssim g(n)$ if there exists $n_{0}$ and a constant $C>0$ such that for all $n\geq n_{0}$, $f(n)\leq Cg(n)$. We also write $f(n)\ll g(n)$ to indicate that $f$ is asymptotically dominated by $g$, namely, for every fixed real number $C>0$, there exists $n(C)$ such that $g(n)\geq Cf(n)$ for all $n>n(C)$. For any positive integer $m\geq 1$, we denote the set $\{1,\ldots,m\}$ by $[m]$. Let $\bm{e}_{i}$ denote the $i^{\text{th}}$ canonical vector, namely, the vector of all zeros except a one in the $i^{\text{th}}$ coordinate. We denote the $\ell_{2}$ norm of a vector $\bm{v}$ by $\left\|\bm{v}\right\|_{\ell_{2}}$, and the operator norm of a matrix ${\bm{M}}$ by $\left\|{\bm{M}}\right\|$. The unit $d$-dimensional sphere (i.e., the set of $d$-dimensional vectors with unit norm) is denoted by $\mathbb{S}^{d-1}$. For a random variable $X$, its sub-gaussian norm, denoted by $\|X\|_{\psi_{2}}$, is defined as $\|X\|_{\psi_{2}}:=\sup_{q\geq 1}q^{-1/2}\big{(}\operatorname{\mathbb{E}}|X|^{q}\big{)}^{1/q}$. [^2] We refer to (Vershynin,, 2010, Lemma 5.5) for other equivalent definitions. In addition, for random vector ${\bm{X}}\in\mathbb{R}^{d}$, its sub-gaussian norm is defined as $\|{\bm{X}}\|_{\psi_{2}}=\sup_{{\bm{z}}\in\mathbb{S}^{d-1}}\|\langle{\bm{X}},{\bm{z}}\rangle\|_{\psi_{2}}$. We denote the CDF and complementary CDF (CCDF) of a standard normal variable (i.e., distributed as ${\mathcal{N}}(0,1)$) by $\Phi(.)$ and $\Phi^{\text{c}}(.)$, respectively.
Definition 3.1 (Label Differential Privacy (DP))
@@
Local label DP training involves injecting noise into the labels and then training with these noisy labels. The simplest way of injecting label noise to ensure label DP is randomized response (RR) introduced by Warner, (1965). Specifically, suppose we require $\epsilon$-labelDP for a problem with $C$ classes, then the distribution of the output $\widehat{y}$ of RR when the true label is $y$ is as follows:
$\mathbb{P}\big{(}\widehat{y}=z\big{)}=\begin{cases}\frac{e^{\epsilon}}{e^{\epsilon}+C-1}\quad\text{ for }z=y,\\ \frac{1}{e^{\epsilon}+C-1}\quad\text{ otherwise}.\end{cases}$
Our label noise model in Section 4 (eq. 4.2) is actually RR for 2 classes. Based on vanilla RR, more sophisticated ways to inject label noise for better performance under label DP have been proposed (Ghazi et al.,, 2021; Malek Esmaeili et al.,, 2021). Our empirical results in Section 5 are with RR and the method of Ghazi et al., (2021).
@@ 4 Full Retraining in the Presence of Label Noise: Theoretical Analysis
Here we will analyze full retraining (as introduced in Section 1) for a linear setting with noisy labels. Since full retraining is the only kind of retraining we consider here, we will omit the word “full” subsequently in this section.
Problem Setting. We consider binary classification under a discriminative mixtures of Gaussian data model with a positive margin. We will first describe the classical Gaussian mixture model and then propose a new model which is endowed with a positive margin.
In the classical Gaussian mixture model, each data point belongs to one of two classes {±$\{\pm 1\}$ with corresponding probabilities $\pi_{+}$, $\pi_{-}$, such that $\pi_{+}+\pi_{-}=1$. Denoting by $y_{i}\in\{-1,+1\}$ the label for data point $i$, the feature vectors $\bm{x}_{i}\in\mathbb{R}^{d}$, for $i\in[n]$, are generated independently as $\bm{x}_{i}\sim{\mathcal{N}}(y_{i}{\bm{{\mu}}},\bm{\Sigma})$, where ${\bm{{\mu}}}\in\mathbb{R}^{d}$ and $\bm{\Sigma}\in\mathbb{R}^{d\times d}$. In other words the mean of feature vectors are ±$\pm{\bm{{\mu}}}$ depending on its class, and $\bm{\Sigma}$ is the covariance of the features. We let $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$.
Gaussian mixture model with positive margin: In this model, each data point $(\bm{x},y)$ is generated independently by first sampling the label $y\in\{-1,+1\}$ with the corresponding probabilities $\pi_{-},\pi_{+}$, and then generating the feature vector $\bm{x}\in\mathbb{R}^{d}$ as:
$\bm{x}=y(1+u){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\bm{z}\,,\text{ where }$ (4.1)
  • $u$ is drawn independently for each sample from a common sub-gaussian distribution with positive support ($u>0$).
  • $\bm{z}\sim{\mathcal{N}}(\bm{0},\bm{I}_{d})$ independent of $u$.
  • $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ is the covariance matrix in the space orthogonal to ${\bm{{\mu}}}$. We have $\bm{\Sigma}^{1/2}{\bm{{\mu}}}=\bm{0}$ and $\bm{\Sigma}$ is of rank $d-1$.
We let $\bm{\Sigma}$ be an arbitrary positive semi-definite matrix to handle general covariances. Let $\lambda_{\min}>0$ and $\lambda_{\max}$ denote the minimum non-zero eigenvalue and the maximum eigenvalue of $\bm{\Sigma}$, respectively. Also without loss of generality, we assume that $u$ has unit sub-gaussian norm ($\|u\|_{\psi_{2}}=1$).
Note that under this model, the projection of datapoints on the orthogonal space of ${\bm{{\mu}}}$ is distributed as normal vectors. On the direction ${\bm{{\mu}}}$, we have $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}$, and so the randomness of data along this direction is modeled in $u$. Since $u$ has a positive support, $\text{sign}(\langle\bm{x},{\bm{{\mu}}}\rangle)=y$ and therefore ${\bm{{\mu}}}$ is a separator of the data (based on its labels). In addition, we have a margin of at least $\gamma=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}$; this is because $\frac{y\langle\bm{x},{\bm{{\mu}}}\rangle}{\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}}=(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}\geq\gamma$. We will mostly refer to the margin $\gamma$ as the “degree of separation”.
We are given access to a training set $\mathcal{T}=\{(\bm{x}_{i},\widehat{y}_{i})\}_{i\in[n]}$ where for each $i\in[n]$, $\widehat{y}_{i}$ is a noisy version of the true label $y_{i}$ (which we do not observe). Specifically:
$\widehat{y}_{i}=\begin{cases}y_{i}&\text{with probability }1-p\,,\\ -y_{i}&\text{with probability }p\,,\end{cases}$ (4.2)
for some $p<1/2$, and independently for each $i\in[n]$. [^3] In the context of $\epsilon$-labelDP for a binary classification problem with randomized response, $p=\frac{1}{e^{\epsilon}+1}$.
4.1 Vanilla Training
Given the training set $\mathcal{T}=\{(\bm{x}_{i},\widehat{y}_{i})\}_{i\in[n]}$, we consider the following linear classifier (Carmon et al.,, 2019):
${\widehat{\bm{\theta}}}_{0}=\frac{1}{n}\sum_{i=1}^{n}\widehat{y}_{i}\bm{x}_{i}\,.$ (4.3)
This classifier’s predicted label for a sample $\bm{x}\in\mathbb{R}^{d}$ is ${\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}_{0}\rangle)$. Note that $\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle=(1/n)\sum_{i}\widehat{y}_{i}\langle\bm{x}_{i},\bm{x}\rangle$ is an average of noisy labels in the training set with weights given by $\langle\bm{x}_{i},\bm{x}\rangle/n$. Therefore, it is similar to kernel methods with the inner product kernel.
Our next result bounds the probability of ${\widehat{\bm{\theta}}}_{0}$ correctly classifying a given fixed test point $\bm{x}$.
Theorem 4.1 (Vanilla Training)
Consider $\bm{x}\notin\mathcal{T}$ and let $y$ be its true label. We have
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)$ $\leq\alpha_{0}(\bm{x}):=1-\frac{1}{2\sqrt{2\pi}}\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\,\text{ and }$
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)=y)$ $\geq\widetilde{\alpha}_{0}(\bm{x}):=1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
The proof of Theorem 4.1 is in Appendix B. Notice that the learned classifier $\widehat{\bm{\theta}}_{0}$ is more likely to be wrong on the samples that are less aligned with (or closer to orthogonal to) the ground truth separator, i.e., ${\bm{{\mu}}}$. We can view $\widehat{\bm{\theta}}_{0}$ as a noisy label provider (a.k.a. pseudo-labeler) where the degree of label noise is non-uniform or sample-dependent unlike the original noisy source used to learn $\widehat{\bm{\theta}}_{0}$. Specifically, for a sample $\bm{x}$ with true label $y$ and predicted label $\widetilde{y}={\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)$, we have: